# AC Josephson Signatures of the Superconducting Higgs Mode

### v5

@@ Abstract
Model.– We consider $s$-wave BCS SC Hamiltonians [71, 72, 73, 74]. The device consists of two SC leads of length $L$ satisfying $\lambda_{F}\ll L\sim\xi_{\text{sc}}$ where $\lambda_{F}$ is the Fermi wavelength and $\xi_{\text{sc}}$ is the coherence length. This is because the Higgs modulations occur over a length-scale $\sim\xi_{\text{sc}}$. Since we do not explicitly model electromagnetic screening, our results are, in principle, valid for junctions with widths comparable to the electromagnetic penetration depth. However, the time-dependent phenomena obtained in this work should remain qualitatively unaffected even for wider junctions. The outer ends of the leads are connected to macroscopic SC reservoirs. For computational simplicity, here we focus on a two-dimensional planar junction with specular tunneling. A similar three-dimensional system, assuming a uniform OP along the transverse directions, is analysed in the SM [59]. The results confirm that our conclusions remain unchanged. The left (L) and right (R) leads contain $N_{L}$ and $N_{R}$ sites, respectively, along the longitudinal direction, and $N_{w}$ sites along the transverse direction. Both leads are connected to infinite superconducting reservoirs. The mean-field Hamiltonian is $H=H_{L}+H_{R}+H_{T}$, where,
@@
$H_{T}=$ ${\textstyle\sum\limits_{j_{y},\sigma}}\big{[}-\mathcal{T}e^{i\frac{\phi(t)}{2}}c_{L1,j_{y},\sigma}^{\dagger}c_{R1,j_{y},\sigma}+h.c.\big{]}.$ (4)
@@
where we have used the Einstein summation convention, $G^{r/a}$ is the retarded/advanced Green’s function, $g^{r/a}_{mn}(\omega)=g^{r/a}(\omega+m\Omega)\delta_{mn}$ is defined in the absence of tunneling, and $\Sigma$ is the self-energy which we detail below. Following standard procedure [101], we partition the system into two leads connected at the junction, which we treat exactly, and two semi-infinite reservoirs attached to the far-ends of the leads. The self-energy consists of four terms: (i) tunneling self energy $\Sigma^{r/a}_{T,RL,m-n}=\Sigma^{r/a^{\scriptstyle\ast}}_{T,LR,n-m}=-[V_{m-n}(\tau_{0}+\tau_{3})/2-V_{-m+n}^{*}(\tau_{0}-\tau_{3})/2]$, where $V_{m-n}=(1/T)\int(d\omega/(2\pi))e^{i(m-n)\Omega t}\mathcal{T}e^{-i\phi(t)/2}$, with $\tau_{\mu}$ denoting the Pauli matrices in Nambu space, (ii) OP self energy, $\Sigma^{r/a}_{\delta\Delta,m-n}=[\Delta_{m-n}\tau_{+}+\Delta_{n-m}^{*}\tau_{-}]$ for $m\neq n$, where $\Delta_{m-n\neq 0}=(1/T)\int(d\omega/(2\pi))e^{i(m-n)\Omega t}\delta\Delta(t)$, (iii) reservoir self-energy, $\Sigma^{r/a/<}_{\text{res.},m-n}=\zeta^{2}\tau_{3}g_{b,m-n}\tau_{3}$, where $g_{b}$ is the boundary Green’s function [101], acting only on the lead sites immediately neighbouring the reservoir ($L1/RN_{R}$ ), (iv) broadening self-energy $\Sigma^{r/a}_{\Gamma}=\pm i\Gamma/2$ and $\Sigma^{<}_{\Gamma}=-i\Gamma f(\omega)$, where $f(\omega)$ is the Fermi function. It aids numerical convergence, and accounts for the lifetime arising from, e.g., relaxation to the quasiparticle continuum, electron-phonon interaction, etc [69]. Following this prescription, the Floquet components of the current are obtained as [69, 76, 93],
$\begin{split}I_{n}=&\sum_{l,m}e\int_{0}^{\Omega}\frac{d\omega}{2\pi}\mathbf{tr}\big{[}\tau_{3}\Sigma_{T,LR,m+n-l}G_{R1LN_{1},lm}^{<}(\omega)\\ &-(L\leftrightarrow R)\big{]}.\end{split}$ (8)
@@
Finally, using a perturbative approach, we show that the Keldysh-Gorkov formulation reduces to the phenomenological model presented earlier, establishing the qualitative equivalence. Starting with the bulk equilibrium gap $\Delta_{0}$, we write $\Delta(t,x)=\Delta_{0}+\delta\Delta(t,x)$ for each lead. The corresponding self-energy is $\Sigma_{\delta\Delta}(t,x)=\delta\Delta(t,x)\tau_{x}$. The tunneling self energy is $\Sigma_{T,LR}(t)=\Sigma_{T,RL}^{*}(t)=(-\mathcal{T})\tau_{z}\exp(-i\phi(t)\tau_{z}/2)$. As detailed in the SM [59], expanding the lesser Green’s functions $G_{L/R}^{<}$ up to $\mathcal{O}(\mathcal{T}^{3})$, $G_{L/R}^{<}=g_{L/R}^{<}+\sum_{j=1}^{3}(\delta G_{L/R}^{<})^{(j)}$, we obtain Eq. (2) with
@@
Here, $\delta(G_{L/R}^{<})^{(1)}$ has one instance of $\Sigma_{\delta\Delta}$, and accounts for the intrinsic dynamics of the pairs. It defines the Higgs susceptibility. $\delta(G_{L/R}^{<})^{(2)}$ has two instances of $\Sigma_{T}$. It describes the effect of the radiating pairs tunneling across the junction, and defines the source term. Finally, $\delta(G_{L/R}^{<})^{(3)}$ has one instance of $\Sigma_{\delta\Delta}$ and two instances of $\Sigma_{T}$. It captures the cross-coupling between the OP oscillations of the two leads mediated by the tunneling pairs. We relegate the expressions for Eq. (9) to the SM [59], focussing only on their physical behaviour here.

### v6

@@ Abstract
Model.– We consider $s$-wave BCS SC Hamiltonians [71, 72, 73, 74]. The device consists of two SC leads of length $L$ satisfying $\lambda_{F}\ll L\sim\xi_{\text{sc}}$ where $\lambda_{F}$ is the Fermi wavelength and $\xi_{\text{sc}}$ is the coherence length. This is because the Higgs modulations occur over a length-scale $\sim\xi_{\text{sc}}$. Since we do not explicitly model electromagnetic screening, our results are, in principle, valid for junctions with widths comparable to the electromagnetic penetration depth. However, the time-dependent phenomena obtained in this work should remain qualitatively unaffected even for wider junctions. The outer ends of the leads are connected to macroscopic SC reservoirs. For computational simplicity, here we focus on a two-dimensional planar junction with specular tunneling. A similar three-dimensional system, assuming a uniform OP along the transverse directions, is analysed in the SM [59]. The results confirm that our conclusions remain unchanged. The left (L) and right (R) leads contain $N_{L}$ and $N_{R}$ sites, respectively, along the longitudinal direction, and $N_{w}$ sites along the transverse direction. Both leads are connected to infinite superconducting reservoirs. The mean-field Hamiltonian is $H=H_{L}+H_{R}+H_{\mathcal{T}}$, where,
@@
$H_{\mathcal{T}}=$ ${\textstyle\sum\limits_{j_{y},\sigma}}\big[-\mathcal{T}e^{i\frac{\phi(t)}{2}}c_{L1,j_{y},\sigma}^{\dagger}c_{R1,j_{y},\sigma}+h.c.\big].$ (4)
@@
where we have used the Einstein summation convention, $G^{r/a}$ is the retarded/advanced Green’s function, $g^{r/a}_{mn}(\omega)=g^{r/a}(\omega+m\Omega)\delta_{mn}$ is defined in the absence of tunneling, and $\Sigma$ is the self-energy which we detail below. Following standard procedure [101], we partition the system into two leads connected at the junction, which we treat exactly, and two semi-infinite reservoirs attached to the far-ends of the leads. The self-energy consists of four terms: (i) tunneling self energy $\Sigma^{r/a}_{\mathcal{T},RL,m-n}=\Sigma^{r/a^{\scriptstyle\ast}}_{\mathcal{T},LR,n-m}=-[V_{m-n}(\tau_{0}+\tau_{3})/2-V_{-m+n}^{*}(\tau_{0}-\tau_{3})/2]$, where $V_{m-n}=(1/T)\int(d\omega/(2\pi))e^{i(m-n)\Omega t}\mathcal{T}e^{-i\phi(t)/2}$, with $\tau_{\mu}$ denoting the Pauli matrices in Nambu space, (ii) OP self energy, $\Sigma^{r/a}_{\delta\Delta,m-n}=[\Delta_{m-n}\tau_{+}+\Delta_{n-m}^{*}\tau_{-}]$ for $m\neq n$, where $\Delta_{m-n\neq 0}=(1/T)\int(d\omega/(2\pi))e^{i(m-n)\Omega t}\delta\Delta(t)$, (iii) reservoir self-energy, $\Sigma^{r/a/<}_{\text{res.},m-n}=\zeta^{2}\tau_{3}g_{b,m-n}\tau_{3}$, where $g_{b}$ is the boundary Green’s function [101], acting only on the lead sites immediately neighbouring the reservoir ($L1/RN_{R}$ ), (iv) broadening self-energy $\Sigma^{r/a}_{\Gamma}=\pm i\Gamma/2$ and $\Sigma^{<}_{\Gamma}=-i\Gamma f(\omega)$, where $f(\omega)$ is the Fermi function. It aids numerical convergence, and accounts for the lifetime arising from, e.g., relaxation to the quasiparticle continuum, electron-phonon interaction, etc [69]. Following this prescription, the Floquet components of the current are obtained as [69, 76, 93],
$\begin{split}I_{n}=&\sum_{l,m}e\int_{0}^{\Omega}\frac{d\omega}{2\pi}\mathbf{tr}\big[\tau_{3}\Sigma_{\mathcal{T},LR,m+n-l}G_{R1LN_{1},lm}^{<}(\omega)\\ &-(L\leftrightarrow R)\big].\end{split}$ (8)
@@
Finally, using a perturbative approach, we show that the Keldysh-Gorkov formulation reduces to the phenomenological model presented earlier, establishing the qualitative equivalence. Starting with the bulk equilibrium gap $\Delta_{0}$, we write $\Delta(t,x)=\Delta_{0}+\delta\Delta(t,x)$ for each lead. The corresponding self-energy is $\Sigma_{\delta\Delta}(t,x)=\delta\Delta(t,x)\tau_{x}$. The tunneling self energy is $\Sigma_{\mathcal{T},LR}(t)=\Sigma_{\mathcal{T},RL}^{*}(t)=(-\mathcal{T})\tau_{z}\exp(-i\phi(t)\tau_{z}/2)$. As detailed in the SM [59], expanding the lesser Green’s functions $G_{L/R}^{<}$ up to $\mathcal{O}(\mathcal{T}^{3})$, $G_{L/R}^{<}=g_{L/R}^{<}+\sum_{j=1}^{3}(\delta G_{L/R}^{<})^{(j)}$, we obtain Eq. (2) with
@@
Here, $\delta(G_{L/R}^{<})^{(1)}$ has one instance of $\Sigma_{\delta\Delta}$, and accounts for the intrinsic dynamics of the pairs. It defines the Higgs susceptibility. $\delta(G_{L/R}^{<})^{(2)}$ has two instances of $\Sigma_{\mathcal{T}}$. It describes the effect of the radiating pairs tunneling across the junction, and defines the source term. Finally, $\delta(G_{L/R}^{<})^{(3)}$ has one instance of $\Sigma_{\delta\Delta}$ and two instances of $\Sigma_{\mathcal{T}}$. It captures the cross-coupling between the OP oscillations of the two leads mediated by the tunneling pairs. We relegate the expressions for Eq. (9) to the SM [59], focussing only on their physical behaviour here.